OpenStudy (anonymous):

Solve the inequality 4x − 6 > 6x − 20. x > 7 x < 7 x < 2 x > 2

OpenStudy (decentnabeel):

4x − 6 > 6x − 20 add 6 on both sides 4x-6+6>6x-20+6 4x>6x-14 subtract6x from both sides 4x-6x>6x-6x-14 -2x>-14 \[Multiply\:both\:sides\:by\:-1\:\left(reverse\:the\:inequality\right)\] \[\left(-1\right)\left(-2x\right)<\left(-1\right)\left(-14\right)\] -2x<-14 divided 2 on both sides -2x/-2<-14/-2 x<7

OpenStudy (decentnabeel):

second option is the correct

OpenStudy (decentnabeel):

@surjithayer ... 2x<14 x<2 this is wrong

OpenStudy (anonymous):

2x<14 x<7

OpenStudy (anonymous):

thank you i was so lost ... now i can take notes because i didnt know where to start thanks again :)

OpenStudy (decentnabeel):

no problem @_lexx.